OpenStudy (anonymous):

Need help w/ linear differential equations... If: L1=D+1, L2=D-2X^2 ; where D is your differential operator. What is L1L2 and L2L1?

OpenStudy (anonymous):

what is L1 & L2...?

OpenStudy (anonymous):

function,operator...?

OpenStudy (anonymous):

Its in the question, D is the differential operator.

OpenStudy (anonymous):

if I'm reading it correctly...: L1L2=(D+1)*(D-2x^2)=D^2 -D(2x^2) +D(1) - 2x^2=D^2 -4x -2x^2

OpenStudy (anonymous):

The answer for L1L2 is: D^2+(D-2x)D-2x(x+2)

OpenStudy (anonymous):

I just don't understand how they got that..

OpenStudy (anonymous):

me too... looking at that now... may be light will come up..:) give me a min

OpenStudy (anonymous):

o.k thanks.

OpenStudy (anonymous):

what they are asking is to show that L1L2 does not equal L2L1.

OpenStudy (anonymous):

can't get it... I see D^2 + (D-2x)D... but where is -2x(x+2) comes from?

OpenStudy (anonymous):

yeah, thats what I thought...

OpenStudy (anonymous):

I dont know if they are doing some product rule....

OpenStudy (anonymous):

L2L1 is D^2+(1-2X)D-2x^2

OpenStudy (anonymous):

I'll play with it... let you know

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

still getting L1L2=L2L1... did you try internet?

OpenStudy (anonymous):

yeah. I dont think that they are just multiplying them together